Solve the equation for the indicated variable. v=1/2 at^2; for t
Answer by stanbon(75887)

Solve the equation for the indicated variable. v=1/2 at^2; for t
---------
v = (1/2) at^2
2t = at^2
t^2 = (2t)/a
t = sqrt[(2t)/a)]
=========================
Cheers,
Stan H.